a block of mass m is held against a rough vertical wall by pressing in with a finger if u is the coeffcieng of friction between block and wall and g is the acceleration due to gravity calculate the minimum force applied by finger to hold the block against wall

Grade:11

2 Answers

Vikas TU
7 years ago
This time normal force = m*a
f =  u*ma or = u*mg that would be the only force would be left for balancing the forec applied by the finger to the block against the wall.
Nikhil Negi
7 years ago
To hold the block against the wall, sum of all external forces must result zero (for block to be at rest w.r.t wall)As the block is pushed against the wall, block would have a tendency to slip downward..So, friction would act up along wall to stop relative motion... Therefore for block to be at equilibrium,f=mg where f=friction forceuR=mg eq1 Where R is reaction contact force by wallApplied force=RUsing eq1 R=mg÷uHence, applied force must be= mg/u